# Beyond Pipelines: A Survey of the Paradigm Shift toward Model-Native Agentic AI

### v1

@@ 5.5. Summary and Discussion
Short-term memory was initially developed to compensate for models’ limited context processing abilities. Early pipeline-based approaches, such as MemoChat (Fei, 2023), used structured templates and decomposable modules to filter and arrange evidence before it reached the model. While interpretable and easy to implement, these systems were limited by cross-module error accumulation and inflexibility, casting the model as a passive participant. As native context windows have expanded, the focus has shifted from mere capacity to effective context management. The challenge is no longer how much information can be processed, but how well the model can filter noise, locate critical evidence, and utilize it for complex reasoning. This has driven the evolution from coarse-grained pipelines to more intelligent, learned policies.
@@
This transition is not an abrupt replacement but a gradual process of internalizing validated pipeline functions into the model’s native abilities. A key bottleneck in this path lies in long-sequence data synthesis and curriculum design. Future breakthroughs will likely involve training models to endogenously learn operations like retrieval, compression, and verification as explicit objectives (Borgeaud et al., 2021; Izacard et al., 2022; Zhao et al., 2025b). As models learn attention mechanisms akin to a retriever and resource allocation strategies similar to a router (Fang et al., 2025), external modules like vector stores will recede from online decision-making. Their role will shift to serving as backend infrastructure for compliance, persistent storage, and offline content preparation, supporting the model’s native governance capabilities.
@@
Long-term memory began primarily as external repositories, with engineering practices emphasizing traceability, control, and governance. Retrieval relied on cascaded re-ranking (Karpukhin et al., 2020), but this design faced two key limitations: the content was not rewritable online, and there was often a mismatch between retrieved relevance and actual utility in the reasoning chain (Asai et al., 2023). A key recent trend is personalization, where long-term memory evolves from a unified knowledge warehouse into a dynamic, user-specific asset composed of traceable evidence cards (Park et al., 2023; Packer et al., 2023). A second trend is the deepening coupling with short-term memory. Although most long-term stores are not yet fully model-native due to their offline nature, methods are emerging that blur this boundary. This includes co-training the retriever and generator to improve evidence utilization (Izacard et al., 2022; Zhang et al., 2024c), using online updates to write back critical facts, and employing lightweight adapters to inject knowledge into model parameters (Wang et al., 2021). Systems like Mem0 (Chhikara et al., 2025) sit at this intersection, combining the governability of external memory with the fast read-write capabilities needed for session-level adaptation, pointing toward an eventual unification of long-term and short-term memory systems.
@@ 6.1.1. Overview
In contrast, the core idea of model-native methods is to internalize search strategies into model parameters through reinforcement learning, enabling end-to-end autonomous research capability. This paradigm enhances long-horizon reasoning consistency and allows the model to dynamically adapt to diverse search contexts. From the perspective of reasoning consistency, internalizing strategies at the parametric level leverages stored knowledge to preserve goal and evidence coherence across multiple turns (Song et al., 2025b), thereby improving reasoning stability and reducing phenomena such as goal drift, lost-in-conversation, and information forgetting in extended prompt chains (Laban et al., 2025; Liu et al., 2023b; Wang et al., 2025b). With respect to adaptability, the model dynamically adjusts retrieval depth and evidence integration according to task complexity, without relying on predefined workflows, thereby reducing unnecessary tool calls, redundant information retrieval, and overall token costs (Jin et al., 2025). By shifting from external orchestration to internalized capability, this paradigm ensures stable reasoning across diverse environments and task structures, enhancing the generalization of Deep Research agents to open-domain, multi-hop, and unstructured evidence tasks.
@@ 6.1.2. Pipeline-based Paradigm
  • Query enhancement.  Query enhancement improves retrieval effectiveness by rewriting, expanding, or constructing task-specific queries (Wang et al., 2023a), thereby increasing retrieval quality and answerability. Query2doc (Wang et al., 2023b) prompts an LLM in a few-shot manner to generate pseudo-documents, which are concatenated with the original query for search, effectively expanding query scope. Ma et al. (2023) introduce a query rewriting framework for RAG, where an LLM or a lightweight model first rewrites the query, retrieves results via the Bing API, and then passes the retrieved content to a reader model for final generation. FreshLLMs (Vu et al., 2023) focus on time-sensitive knowledge by combining multiple evidence sources with structured metadata, and guide attention according to temporal order. Jin et al. (2024) present BIDER, which improves retrieval coverage and matching accuracy through bidirectional expansion of both documents and queries from the index side.
  • Evidence processing.  To ensure that retrieved information is relevant, concise, and structurally coherent before being injected into the model context, evidence processing refines and organizes raw retrieval outputs. Xu et al. (2023) propose RECOMP, which introduces a compression layer prior to injection, combining extractive and abstractive compressors that can selectively omit uninformative content. Jiang et al. (2023b) develop LongLLMLingua, applying global reranking and span-level extraction to mitigate context explosion. Vu et al. (2023) normalize search results into structured evidence lists for more consistent integration. Li et al. (2024b) design CorpusLM, which treats a unified corpus as fixed external memory, retrieving fragments online and directly reading from a clean repository during inference. Li et al. (2024d) present RetroLLM, converting web pages into Markdown to build writable memory, enabling continuous accumulation and reuse of retrieved evidence. Edge et al. (2024a) proposed GraphRAG that replace scattered documents with knowledge graphs and community summaries, supporting path-level interpretability and global evidence aggregation.
@@ 6.1.3. Model-native Paradigm
Offline training conducts external information retrieval within closed, locally hosted databases that operate without internet access. This setup provides a stable and controllable training environment, allowing the model to learn retrieval and reasoning behaviors under low noise and fixed data conditions. Representative works such as Search-R1 (Jin et al., 2025), R1-Searcher (Song et al., 2025a), ReSearch (Chen et al., 2025b), R-Search (Zhao et al., 2025a), and R1-Searcher++ (Song et al., 2025b) adopt Wikipedia as their static retrieval corpus, enabling supervised or reinforcement learning over consistent textual contexts. Similarly, M2IO-R1 (Xiao et al., 2025b) employs a pre-constructed dataset (M2IO-Inserter) to simulate information insertion and retrieval tasks in a controlled manner. Such designs ensure reproducibility and avoid issues commonly encountered in online environments, including network instability, API rate limitations, and unpredictable web responses.
@@ 6.1.4. Summary and Discussion
    Another line of research prioritizes reproducibility and training stability by using static databases, such as a Wikipedia dump, as the external knowledge source. This controlled environment allows for more rigorous evaluation of an agent’s reasoning and retrieval strategies without the noise and non-stationarity of the open web. Representative works in this category include Search-R1 (Jin et al., 2025), ReSearch (Chen et al., 2025b), R-Search (Zhao et al., 2025a), and the R1-Searcher series (Song et al., 2025a, b). M2IO-R1 (Xiao et al., 2025b) further adapts this approach for multimodal scenarios by retrieving from a pre-constructed database of text-image pairs. To bridge the gap between static and live environments, ZeroSearch (Sun et al., 2025b) introduces a novel training strategy where an LLM is used to simulate a search engine, generating both relevant and noisy documents. This allows the agent to be trained with curriculum-based difficulty, improving robustness while avoiding the high cost and instability of live API calls.
@@
  • LLM.  Search-R1 (Jin et al., 2025), R1-Searcher (Song et al., 2025a), ReSearch (Chen et al., 2025b), R-Search (Zhao et al., 2025a) utilize a single LLM that, after reinforcement learning for deep research work, becomes an LRM that internalizes tool usage mastery, responsible for completing the entire process of reasoning, generating search queries, understanding retrieval results, and generating answers.
@@
  • Process rewards.  Process rewards provide stepwise feedback, guiding models to balance reasoning depth, efficiency, and traceability. R1-Searcher++ (Song et al., 2025b) combines format, outcome, and group rewards, encouraging accurate answers with minimal retrievals. R-Search (Zhao et al., 2025a) extends this approach with multi-stage reward mechanisms to jointly optimize reasoning quality and resource efficiency.
@@ 6.2.2. Pipeline-based Paradigm
  • Reflective agents.  This approach introduces a self-evaluation and correction loop, enabling the agent to identify and rectify its own errors. The Reflexion (Shinn et al., 2023) framework allows an agent to analyze task failures and store these reflections in memory to avoid repeating mistakes, which has been shown to improve success rates. Similarly, the RCI (Kim et al., 2023) framework demonstrated increased accuracy in GUI form-filling tasks by incorporating a self-correction mechanism. In mobile scenarios, Mobile-Agent-V (Wang et al., 2025h) achieves self-correction through a fusion of visual backtracking and linguistic reflection, reviewing past demonstration videos to evaluate its actions and generate improvement strategies.
@@ 6.2.3. Model-native Paradigm
    Research in planning module follows two main approaches: imitative and generative planning. Imitative planning learns from expert demonstrations. For instance, CoAT(Zhang et al., 2024e) makes the “chain of action” and “chain of thought” explicit for supervised fine-tuning, enhancing interpretability and robustness. Similarly, WebAgent(Gur et al., 2024) constructs a modular pipeline to normalize and orchestrate natural language instructions. In contrast, generative planning enables agents to create plans more autonomously. WMA-Agent(Chae et al., 2024), for example, allows the planner to “imagine before deciding” by learning the environment’s dynamics, showing performance gains on benchmarks like WebArena. Another approach, WebSynthesis(Gao et al., 2025b), uses a world model with MCTS to synthesize high-quality training trajectories, thereby augmenting its policy learning process.
@@
    In discriminative reflection, the focus is on step-by-step verification. For instance, STEVE(Lu et al., 2025e) reflects on each action’s validity by using a large multimodal model to verify the screen transition (previous screen, action, next screen). V-Droid(Dai et al., 2025), on the other hand, reflects by using a verifier to evaluate a set of candidate actions based on learned task progress preferences. In evaluative reflection, the goal is to assess the overall quality of a sequence. GUI-PRA(Xiong et al., 2025) employs a Process Reward Model (PRM) to reflect on entire action sequences, using this evaluation to shape rewards during training and re-rank actions during inference. Similarly, UI-Genie(Xiao et al., 2025a) uses a learned reward model to score actions. This reflective scoring serves a dual purpose: re-ranking actions during inference and providing a reward signal for fine-tuning, thereby creating a self-improvement loop.
@@
    The earliest work in this area, UI-TARS(Qin et al., 2025), was the first to implement end-to-end online reinforcement learning in a native application environment. It demonstrated the online learnability of a “pixels-to-actions” policy by having a model take screenshots as direct input and produce action commands as output, using real interactions and environmental feedback to optimize its policy function. Subsequently, ARPO(Lu et al., 2025d) built upon this foundation by proposing a mechanism that decouples experience replay from policy updates. By converting collected GUI trajectories into reusable samples and introducing a high-quality task filtering strategy, it achieved efficient and stable convergence for end-to-end reinforcement learning. This allowed the model to continuously improve its perception, planning, and execution capabilities through sustained online interaction. ZeroGUI(Yang et al., 2025b) further advanced this line of research by breaking the dependence on human supervision. It achieved a “zero human cost” online end-to-end training pipeline through automatic task generation, automatic reward signal evaluation, and an adaptive sampling mechanism. This enabled the model to self-explore and correct its policy within a closed loop. In parallel, DART(Li et al., 2025d) significantly improved the efficiency and stability of online training in multi-task, long-sequence environments by introducing a decoupled multi-turn reinforcement learning framework that separates environment exploration from policy updates. Building on these advances, UI-TARS-2(Wang et al., 2025j) extended the training to multi-turn interactions and task transfer. It constructed a complete multi-task online optimization loop, which enabled the agent to achieve policy generalization and self-evolution over long-term interactions. More recent works, such as UItron(Zeng et al., 2025a), Mobile-Agent-v3/GUI-Owl(Ye et al., 2025c) and OpenCUA (Wang et al., 2025e), represent a culmination of this research direction. Both approaches combine reinforcement learning with large multimodal models, allowing a single architecture to handle the three core functions of perception, planning, and execution. By continuously fine-tuning with online reward signals, these agents achieve adaptive, end-to-end capabilities across different platforms and tasks.
@@ 7.1.1. Multi-agent Collaboration
More recent pipeline-based methods have begun to explore the algorithmic generation of workflows to create more dynamic team structures. AFLOW (Zhang et al., 2024f), for example, formulates workflow optimization as a search problem, applying Monte-Carlo Tree Search to explore and refine candidate collaboration structures. Such approaches lay the methodological groundwork for treating the system’s architecture itself as an action space, setting the stage for the more adaptive strategies found in the model-native paradigm.
@@
Early efforts in this paradigm focused on improving the core reasoning and collaborative capabilities of the agents. For instance, MALT (Motwani et al., 2024) uses an offline multi-agent RL approach to improve reasoning by decomposing tasks into distinct generator, verifier, and refiner roles. CORY (Ma et al., 2024a) employs a two-role replication strategy within a cooperative MARL framework, where a single LLM plays both a “leader” and an “observer”, swapping them during training to improve stability. Building on this, more general frameworks have emerged. MARFT (Liao et al., 2025) and MAGRPO (Liu et al., 2025b) adapt algorithms like GRPO to the multi-agent setting to improve training stability without requiring individual value networks, while MATPO (Mo et al., 2025) introduces a principled credit assignment mechanism for training a single model to perform distinct roles.
A more advanced line of research has moved beyond optimizing agent behavior to learning the system’s construction and architecture. Works like MasHost (Yang et al., 2025d), G-Designer (Zhang et al., 2024h), and ARG-Designer (Li et al., 2025g) treat the multi-agent system’s topology—including the number of agents, their roles, and their communication links—as a learnable action space. By using reinforcement learning or graph generation techniques, these methods can autonomously assemble a bespoke multi-agent system tailored to a specific task, elevating the learning objective from behavior to architecture.

### v2

@@ 5.5. Summary and Discussion
Short-term memory was initially developed to compensate for models’ limited context processing abilities. Early pipeline-based approaches, such as MemoChat (Lu et al., 2023a), used structured templates and decomposable modules to filter and arrange evidence before it reached the model. While interpretable and easy to implement, these systems were limited by cross-module error accumulation and inflexibility, casting the model as a passive participant. As native context windows have expanded, the focus has shifted from mere capacity to effective context management. The challenge is no longer how much information can be processed, but how well the model can filter noise, locate critical evidence, and utilize it for complex reasoning. This has driven the evolution from coarse-grained pipelines to more intelligent, learned policies.
@@
This transition is not an abrupt replacement but a gradual process of internalizing validated pipeline functions into the model’s native abilities. A key bottleneck in this path lies in long-sequence data synthesis and curriculum design. Future breakthroughs will likely involve training models to endogenously learn operations like retrieval, compression, and verification as explicit objectives (Borgeaud et al., 2021; Izacard et al., 2022; Zhao et al., 2025c). As models learn attention mechanisms akin to a retriever and resource allocation strategies similar to a router (Liu et al., 2025b), external modules like vector stores will recede from online decision-making. Their role will shift to serving as backend infrastructure for compliance, persistent storage, and offline content preparation, supporting the model’s native governance capabilities.
@@
Long-term memory began primarily as external repositories, with engineering practices emphasizing traceability, control, and governance. Retrieval relied on cascaded re-ranking (Karpukhin et al., 2020), but this design faced two key limitations: the content was not rewritable online, and there was often a mismatch between retrieved relevance and actual utility in the reasoning chain (Asai et al., 2023). A key recent trend is personalization, where long-term memory evolves from a unified knowledge warehouse into a dynamic, user-specific asset composed of traceable evidence cards (Park et al., 2023; Packer et al., 2024). A second trend is the deepening coupling with short-term memory. Although most long-term stores are not yet fully model-native due to their offline nature, methods are emerging that blur this boundary. This includes co-training the retriever and generator to improve evidence utilization (Izacard et al., 2022; Zhang et al., 2024a), using online updates to write back critical facts, and employing lightweight adapters to inject knowledge into model parameters (Wang et al., 2021). Systems like Mem0 (Chhikara et al., 2025) sit at this intersection, combining the governability of external memory with the fast read-write capabilities needed for session-level adaptation, pointing toward an eventual unification of long-term and short-term memory systems.
@@ 6.1.1. Overview
In contrast, the core idea of model-native methods is to internalize search strategies into model parameters through reinforcement learning, enabling end-to-end autonomous research capability. This paradigm enhances long-horizon reasoning consistency and allows the model to dynamically adapt to diverse search contexts. From the perspective of reasoning consistency, internalizing strategies at the parametric level leverages stored knowledge to preserve goal and evidence coherence across multiple turns (Song et al., 2025b), thereby improving reasoning stability and reducing phenomena such as goal drift, lost-in-conversation, and information forgetting in extended prompt chains (Laban et al., 2025; Liu et al., 2023b; Wang et al., 2025a). With respect to adaptability, the model dynamically adjusts retrieval depth and evidence integration according to task complexity, without relying on predefined workflows, thereby reducing unnecessary tool calls, redundant information retrieval, and overall token costs (Jin et al., 2025). By shifting from external orchestration to internalized capability, this paradigm ensures stable reasoning across diverse environments and task structures, enhancing the generalization of Deep Research agents to open-domain, multi-hop, and unstructured evidence tasks.
@@ 6.1.2. Pipeline-based Paradigm
  • Query enhancement.  Query enhancement improves retrieval effectiveness by rewriting, expanding, or constructing task-specific queries (Wang et al., 2023a), thereby increasing retrieval quality and answerability. Query2doc (Wang et al., 2023b) prompts an LLM in a few-shot manner to generate pseudo-documents, which are concatenated with the original query for search, effectively expanding query scope. Ma et al. (2023) introduce a query rewriting framework for RAG, where an LLM or a lightweight model first rewrites the query, retrieves results via the Bing API, and then passes the retrieved content to a reader model for final generation. FreshLLMs (Vu et al., 2023) focus on time-sensitive knowledge by combining multiple evidence sources with structured metadata, and guide attention according to temporal order.
  • Evidence processing.  To ensure that retrieved information is relevant, concise, and structurally coherent before being injected into the model context, evidence processing refines and organizes raw retrieval outputs. Xu et al. (2023) propose RECOMP, which introduces a compression layer prior to injection, combining extractive and abstractive compressors that can selectively omit uninformative content. Jiang et al. (2023b) develop LongLLMLingua, applying global reranking and span-level extraction to mitigate context explosion. Vu et al. (2023) normalize search results into structured evidence lists for more consistent integration. Li et al. (2024b) design CorpusLM, which treats a unified corpus as fixed external memory, retrieving fragments online and directly reading from a clean repository during inference. Edge et al. (2024) proposed GraphRAG that replace scattered documents with knowledge graphs and community summaries, supporting path-level interpretability and global evidence aggregation.Jin et al. (2024) present BIDER, which refines retrieval documents into Key Supporting Evidence (KSE) through knowledge synthesis, supervised fine-tuning, and preference alignment, thereby improving LLM’s answer quality while reducing input length.
@@ 6.1.3. Model-native Paradigm
Offline training conducts external information retrieval within closed, locally hosted databases that operate without internet access. This setup provides a stable and controllable training environment, allowing the model to learn retrieval and reasoning behaviors under low noise and fixed data conditions. Representative works such as Search-R1 (Jin et al., 2025), R1-Searcher (Song et al., 2025a), ReSearch (Chen et al., 2025b), R-Search (Zhao et al., 2025b), and R1-Searcher++ (Song et al., 2025b) adopt Wikipedia as their static retrieval corpus, enabling supervised or reinforcement learning over consistent textual contexts. Similarly, M2IO-R1 (Xiao et al., 2025b) employs a pre-constructed dataset (M2IO-Inserter) to simulate information insertion and retrieval tasks in a controlled manner. Such designs ensure reproducibility and avoid issues commonly encountered in online environments, including network instability, API rate limitations, and unpredictable web responses.
@@ 6.1.4. Summary and Discussion
    Another line of research prioritizes reproducibility and training stability by using static databases, such as a Wikipedia dump, as the external knowledge source. This controlled environment allows for more rigorous evaluation of an agent’s reasoning and retrieval strategies without the noise and non-stationarity of the open web. Representative works in this category include Search-R1 (Jin et al., 2025), ReSearch (Chen et al., 2025b), R-Search (Zhao et al., 2025b), and the R1-Searcher series (Song et al., 2025a, b). M2IO-R1 (Xiao et al., 2025b) further adapts this approach for multimodal scenarios by retrieving from a pre-constructed database of text-image pairs. To bridge the gap between static and live environments, ZeroSearch (Sun et al., 2025b) introduces a novel training strategy where an LLM is used to simulate a search engine, generating both relevant and noisy documents. This allows the agent to be trained with curriculum-based difficulty, improving robustness while avoiding the high cost and instability of live API calls.
@@
  • LLM.  Search-R1 (Jin et al., 2025), R1-Searcher (Song et al., 2025a), ReSearch (Chen et al., 2025b), R-Search (Zhao et al., 2025b) utilize a single LLM that, after reinforcement learning for deep research work, becomes an LRM that internalizes tool usage mastery, responsible for completing the entire process of reasoning, generating search queries, understanding retrieval results, and generating answers.
@@
  • Process rewards.  Process rewards provide stepwise feedback, guiding models to balance reasoning depth, efficiency, and traceability. R1-Searcher++ (Song et al., 2025b) combines format, outcome, and group rewards, encouraging accurate answers with minimal retrievals. R-Search (Zhao et al., 2025b) extends this approach with multi-stage reward mechanisms to jointly optimize reasoning quality and resource efficiency.
@@ 6.2.2. Pipeline-based Paradigm
  • Reflective agents.  This approach introduces a self-evaluation and correction loop, enabling the agent to identify and rectify its own errors. The Reflexion (Shinn et al., 2023) framework allows an agent to analyze task failures and store these reflections in memory to avoid repeating mistakes, which has been shown to improve success rates. Similarly, the RCI (Kim et al., 2023) framework demonstrated increased accuracy in GUI form-filling tasks by incorporating a self-correction mechanism. In mobile scenarios, Mobile-Agent-V (Wang et al., 2025i) achieves self-correction through a fusion of visual backtracking and linguistic reflection, reviewing past demonstration videos to evaluate its actions and generate improvement strategies.
@@ 6.2.3. Model-native Paradigm
    Research in planning module follows two main approaches: imitative and generative planning. Imitative planning learns from expert demonstrations. For instance, CoAT(Zhang et al., 2024c) makes the “chain of action” and “chain of thought” explicit for supervised fine-tuning, enhancing interpretability and robustness. Similarly, WebAgent(Gur et al., 2024) constructs a modular pipeline to normalize and orchestrate natural language instructions. In contrast, generative planning enables agents to create plans more autonomously. WMA-Agent(Chae et al., 2024), for example, allows the planner to “imagine before deciding” by learning the environment’s dynamics, showing performance gains on benchmarks like WebArena. Another approach, WebSynthesis(Gao et al., 2025b), uses a world model with MCTS to synthesize high-quality training trajectories, thereby augmenting its policy learning process.
@@
    In discriminative reflection, the focus is on step-by-step verification. For instance, STEVE(Lu et al., 2025f) reflects on each action’s validity by using a large multimodal model to verify the screen transition (previous screen, action, next screen). V-Droid(Dai et al., 2025), on the other hand, reflects by using a verifier to evaluate a set of candidate actions based on learned task progress preferences. In evaluative reflection, the goal is to assess the overall quality of a sequence. GUI-PRA(Xiong et al., 2025) employs a Process Reward Model (PRM) to reflect on entire action sequences, using this evaluation to shape rewards during training and re-rank actions during inference. Similarly, UI-Genie(Xiao et al., 2025a) uses a learned reward model to score actions. This reflective scoring serves a dual purpose: re-ranking actions during inference and providing a reward signal for fine-tuning, thereby creating a self-improvement loop.
@@
    The earliest work in this area, UI-TARS(Qin et al., 2025), was the first to implement end-to-end online reinforcement learning in a native application environment. It demonstrated the online learnability of a “pixels-to-actions” policy by having a model take screenshots as direct input and produce action commands as output, using real interactions and environmental feedback to optimize its policy function. Subsequently, ARPO(Lu et al., 2025e) built upon this foundation by proposing a mechanism that decouples experience replay from policy updates. By converting collected GUI trajectories into reusable samples and introducing a high-quality task filtering strategy, it achieved efficient and stable convergence for end-to-end reinforcement learning. This allowed the model to continuously improve its perception, planning, and execution capabilities through sustained online interaction. ZeroGUI(Yang et al., 2025b) further advanced this line of research by breaking the dependence on human supervision. It achieved a “zero human cost” online end-to-end training pipeline through automatic task generation, automatic reward signal evaluation, and an adaptive sampling mechanism. This enabled the model to self-explore and correct its policy within a closed loop. In parallel, DART(Li et al., 2025d) significantly improved the efficiency and stability of online training in multi-task, long-sequence environments by introducing a decoupled multi-turn reinforcement learning framework that separates environment exploration from policy updates. Building on these advances, UI-TARS-2(Wang et al., 2025k) extended the training to multi-turn interactions and task transfer. It constructed a complete multi-task online optimization loop, which enabled the agent to achieve policy generalization and self-evolution over long-term interactions. More recent works, such as UItron(Zeng et al., 2025a), Mobile-Agent-v3/GUI-Owl(Ye et al., 2025c) and OpenCUA (Wang et al., 2025e), represent a culmination of this research direction. Both approaches combine reinforcement learning with large multimodal models, allowing a single architecture to handle the three core functions of perception, planning, and execution. By continuously fine-tuning with online reward signals, these agents achieve adaptive, end-to-end capabilities across different platforms and tasks.
@@ 7.1.1. Multi-agent Collaboration
More recent pipeline-based methods have begun to explore the algorithmic generation of workflows to create more dynamic team structures. AFLOW (Zhang et al., 2024d), for example, formulates workflow optimization as a search problem, applying Monte-Carlo Tree Search to explore and refine candidate collaboration structures. Such approaches lay the methodological groundwork for treating the system’s architecture itself as an action space, setting the stage for the more adaptive strategies found in the model-native paradigm.
@@
Early efforts in this paradigm focused on improving the core reasoning and collaborative capabilities of the agents. For instance, MALT (Motwani et al., 2024) uses an offline multi-agent RL approach to improve reasoning by decomposing tasks into distinct generator, verifier, and refiner roles. CORY (Ma et al., 2024a) employs a two-role replication strategy within a cooperative MARL framework, where a single LLM plays both a “leader” and an “observer”, swapping them during training to improve stability. Building on this, more general frameworks have emerged. MARFT (Liao et al., 2025) and MAGRPO (Liu et al., 2025c) adapt algorithms like GRPO to the multi-agent setting to improve training stability without requiring individual value networks, while MATPO (Mo et al., 2025) introduces a principled credit assignment mechanism for training a single model to perform distinct roles.
A more advanced line of research has moved beyond optimizing agent behavior to learning the system’s construction and architecture. Works like MasHost (Yang et al., 2025d), G-Designer (Zhang et al., 2024f), and ARG-Designer (Li et al., 2025g) treat the multi-agent system’s topology—including the number of agents, their roles, and their communication links—as a learnable action space. By using reinforcement learning or graph generation techniques, these methods can autonomously assemble a bespoke multi-agent system tailored to a specific task, elevating the learning objective from behavior to architecture.